# Study of gravitational waves from phase transitions in three-component dark matter

### v2

@@ Abstract
This paper presents a dark matter model comprising three types of particles with distinct spins, along with a scalar field $\phi$ that mediates interactions between Standard Model particles and dark matter. It discusses the electroweak phase transition following the Big Bang, during which all particles are initially massless due to the inactive Higgs mechanism. As temperature decreases, the effective potential reaches zero at two points, leading to two minima at the critical temperature ($T_{c}$), and eventually to a true vacuum state. The formation of new vacuum bubbles, where electroweak symmetry is broken and particles acquire mass, generates gravitational waves as these bubbles interact with the fabric of space-time. The paper derives the gravitational wave frequency and detection range based on the model’s parameters, aligning with observational data from the Planck satellite and detection thresholds from PandaX-4T and XENONnT. It concludes by comparing the predicted background gravitational wave density with the sensitivities of LISA and BBO detectors.
@@ 1 Introduction
After the Big Bang, all particles, including dark matter particles undergoes a electroweak phase transition at a certain temperature. We present a dark matter model composed of three types of particles, each with distinct spins, along with a scalar field $\phi$ that serves as an intermediary between Standard Model particles and dark matter, capable of assuming the vacuum expectation value. Prior to electroweak phase transition, three-component dark matter particles and all SM particles are massless, because the Higgs mechanism has not yet activated. Before the electroweak phase transition, the minimum of effective potential at a certain value of the scalar field $\phi$ is equal to zero, which indicates the initial vacuum before the phase transition. As the temperature decreases, the effective potential becomes zero at two distinct points of the scalar field $\phi$, resulting in two minima corresponding to the critical temperature ($T_{c}$). As the temperature continues to drop, the potential reaches a distinct negative value, leading to a single minimum known as the ”true vacuum”. Cosmic phase transitions happen when the temperature decreases below a critical level, resulting in the Universe transitioning from a symmetric phase to one characterized by broken symmetry. This temperature, at which the cores of true vacuum bubbles begin to form in space, is known as the nucleation temperature ($T_{n}$). In fact, with the breaking of symmetry, the scalar field ($\phi$) acquires a nonzero vacuum expectation value (VEV), and all particles gain mass through the Higgs mechanism. New vacuum bubbles form spherically and move through space, replacing the vacuum before the phase transition. Inside the bubbles, electroweak symmetry is broken, and the Higgs mechanism is activated. In fact, inside the bubbles, dark matter and Standard Model particles have mass, whereas outside these bubbles, they are massless [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
@@
The motion of spherically shaped new vacuum bubbles (true vacuum) within the hot plasma resulting from the Big Bang leads to expansion and oscillation in the space-time fabric due to the distribution of mass in space. When the walls of the bubbles collide with each other, they shake the surrounding space-time, which generates gravitational waves that propagate through space. These gravitational waves were produced in the early universe, and their strength has diminished significantly over time and the expansion of the universe, ultimately transforming into cosmic background gravitational waves. In this paper, we derive the GW frequency and range for the detection of these background gravitational waves, considering the assumption of the existence of multi-component dark matter [11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
@@
In this article, we first present the Lagrangian and its dependent constraints for our dark matter model based on WIMPs. In our model, a pair is considered for the Higgs particle, and its mass ($M_{H_{2}}$) is determined by the Gildener–Weinberg mechanism. Next, we select points from our model’s parameter space where their relic density aligns with the Planck satellite report and their rescaled DM-nucleon cross sections exceed the PandaX-4T and XENONnT detection threshold. We then calculate the effective potential for these points and, using the Euclidean action, determine the necessary parameters for background gravitational waves, which we substitute into the relevant equations. Finally, We determine the detection range for the background gravitational wave density using a three-component dark matter model and compare it to the sensitivities of the LISA and BBO gravitational wave detectors [21, 22, 23, 24, 25, 26, 27, 28].
@@
The paper is organized as follows: Section 2 introduces the Lagrangian of our model, while Section 3 examines the electroweak phase transition via the one-loop effective potential. In Section 4, we present our results, considering the gravitational wave spectrum using two selected points outlined in Table 2. We conclude in Section 5.
@@ 2.1 Lagrangian of model
dark charge ($Q$) $1$ 0 0 $\frac{1}{2}$ $-\frac{1}{2}$ $-\frac{1}{2}$ $\frac{1}{2}$
@@
Consequently, the average $<H_{2}>$ in the flat direction is calculated as follows:
$<H_{2}>=\sqrt{\nu_{1}^{2}+\nu_{2}^{2}}=\nu.$ (2.11)
@@ 2.2 Appropriate points
In this section, we use the micrOMEGAs to select parameter points from our model that achieve a relic density between 0.10 and 0.14, while also ensuring that their DM-nucleon cross section significantly exceeds the detection threshold of the PandaX-4T and XENONnT detectors [30, 21]. For $M_{DM}>40\;GeV$, the direct detection constraint is expressed as follows:
@@
where $\xi_{S}+\xi_{\psi}+\xi_{V}=1$. Using dark matter relic density measurements from the Planck collaboration and considering (2.14), we identify two suitable points presented in Table 2.
Table 2: Characteristics of the selected points with $\lambda_{\phi s}=8$, according to inequality (2.14).
Point $M_{V}$(GeV) $M_{S}$(GeV) $M_{\psi}$(GeV) ${\nu}_{2}$(GeV) $M_{H_{2}}$(GeV) ${\Omega}h^{2}$
A 54.5 1125 611 567 125.79 0.12
B 330 665 175 340 124 0.14
@@
The rescaled DM-nucleon cross sections of selected points, relative to the threshold of the PandaX-4T and XENONnT detectors, are shown in Figure 2.
@@
Figure 2: These figures illustrate the points from table 2 in relation to rescaled DM-Proton cross section versus DM masses, with $\nu_{2}$ and ${\Omega}h^{2}$ represented by color bar.
As shown in figure 2, both points selected fall below the PandaX-4T and XENONnT line threshold. Thus, we can use these points to calculate the gravitational waves from the electroweak phase transition.
@@ 3.1 One-loop effective potential
where $J_{(B_{i})}(x)$ and $J_{(F_{i})}(x)$ are thermal functions defined as follows:
@@
[…]
@@
The nucleation temperature ($T_{n}$) occurs where the corresponding euclidean action is $S_{E}=\frac{S_{3}(T_{N})}{T_{N}}\simeq 140$ [24]. To solve equation (3.12) and find the Euclidean action (3.11), we used the AnyBubble package [23]. This package allows us to calculate the Euclidean action for each selected point using the two minima in the effective potential diagram as function of the scalar field ($\phi$), and plot the ratio of the Euclidean action to temperature (T). In figure 3, we present $\frac{S_{3}(T)}{T}$ for selected points in table 2.
Refer to caption     Refer to caption
@@
Figure 3: The blue line shows $S_{3}(T)/T$ as a function of T, while the dashed red line indicates $S_{3}(T)/T$ = 140, where nucleation occurs. The black dots are derived using the AnyBubble package.
In figure 3, the nucleation temperature ($T_{n}$) is determined by finding the intersection point of the curve fitting and the red line.
@@
To determine the critical temperature $T_{c}$ for the points in table 2, we must plot the effective potential (3.1) against the scalar field ($\phi$) to identify the temperature at which the effective potential exhibits two aligned minima (see figure 4).
Refer to caption     Refer to caption
@@
Figure 4: Effective potential diagrams of points in table 3 as a function of $\phi$ scalar field at nucleation, critical, and supercritical temperatures.
@@
Figure 4 shows the effective potential for each point in Table 2 at the three temperatures: nucleation ($T_{n}$), critical ($T_{c}$), and supercritical ($T>T_{c}$). To ensure $V_{\text{eff}}(0,T)=0$ at all temperatures, we subtract a constant term from the potential: $V_{\text{eff}}(\phi,T)\rightarrow V_{\text{eff}}(\phi,T)-V_{\text{eff}}(0,T)$. As shown in Figure 4, at the nucleation temperature ($T_{n}$), the zero minimum is known as a false vacuum and the negative minimum is known as a true vacuum. There is a bulge between the two minima, which prevents the electroweak phase transition from occurring smoothly between the minima, and instead occurs suddenly and explosively, like a dam breaking.
@@
The supercritical temperatures ($T>T_{c}$) shown in figure 4 are determined as arbitrary and indicate the position of the effective potential in relation to the nucleation and critical temperatures. Figure 4 illustrates that as the temperature decreased, the effective potential also diminishes. Table 3 specifies the calculated values for the nucleation and critical temperatures, as follows:
Table 3: The nucleation and critical temperatures according to points of table 2.
Point $T_{n}$(GeV) $T_{C}$(GeV) ${\phi}_{V_{min}}$(GeV)
A 116 162.5 606.022
B 101 128.2 402.363
@@
In Table 3, the ${\phi}_{V_{min}}$ denotes the point of ${\phi}$ scalar field at which the effective potential value in $T_{n}$ is minimized, as illustrated in Figure 4. In this context, the ${\phi}_{V_{min}}$ represents the true vacuum.
@@ 3.2 Gravitational waves
$\alpha=\frac{\Delta\left(V_{\text{eff}}-T\frac{\partial V_{\text{eff}}}{\partial T}\right)_{\big{|}_{T_{n}}}}{\rho_{*}},\;\;\;\rho_{*}=\frac{\pi^{2}g_{*}}{30}T_{n}^{4}$
$\frac{\beta}{H_{*}}=T_{n}\left.\frac{d}{dT}\left(\frac{S_{3}(T)}{T}\right)\right|_{T_{n}},$ (3.14)
@@
where $g_{*}=100$. In equation (3.2), we obtain the parameter $\frac{\beta}{H_{*}}$ from the slope of the graphs in figure 3 at the intersection point of the curve fitting and the red line. Now, we derive the density relations of gravitational waves for all three types of their sources, including collision (coll), sound wave (sw), and turbulence (turb). We begin with the density of background gravitational waves from collision sources [35]:
@@
$\Omega_{\text{coll}}(f)h^{2}=1.67\times 10^{-5}\left(\frac{\beta}{H_{*}}\right)^{-2}\left(\frac{k\alpha}{1+\alpha}\right)^{2}\left(\frac{g_{*}}{100}\right)^{-\frac{1}{3}}\left(\frac{0.11v_{w}^{3}}{0.42+v_{w}^{2}}\right)S_{\text{coll}},$ (3.15)
where k is the fraction of latent heat converted into the gradient energy of the Higgs-like field, defined as follows:
$k=\frac{1}{1+0.715\alpha}\left(0.715\alpha+\frac{4}{27}\sqrt{\frac{3\alpha}{2}}\right).$ (3.16)
In equation (3.15), $v_{w}$ is the bubble propagation speed, which we take to be equal to one [36]. Additionally, f is the GW frequency of the gravitational wave. $S_{\text{coll}}$ is the spectral shape parameter of the bubble, defined by the following relation:
$S_{\text{coll}}=\frac{3.8\left(\frac{f}{f_{\text{coll}}}\right)^{2.8}}{2.8\left(\frac{f}{f_{\text{coll}}}\right)^{3.8}+1},$
$f_{\text{coll}}=1.65\times 10^{-5}\left(\frac{0.62}{v_{w}^{2}-0.1v_{w}+1.8}\right)\left(\frac{\beta}{H_{*}}\right)\left(\frac{T_{n}}{100}\right)\left(\frac{g_{*}}{100}\right)^{\frac{1}{6}}.$ (3.17)
The density of background GW from sound wave sources is defined as [37]:
$\Omega_{\text{sw}}(f)h^{2}=2.65\times 10^{-6}\left(\frac{\beta}{H_{*}}\right)^{-1}\left(\frac{k_{v}\alpha}{1+\alpha}\right)^{2}\left(\frac{g_{*}}{100}\right)^{-\frac{1}{3}}v_{w}S_{\text{sw}},$ (3.18)
where $k_{v}$ is the bulk motion of the fluid, defined as follows:
$k_{v}=\frac{\alpha}{0.73+0.083\sqrt{\alpha}+\alpha}.$ (3.19)
@@
In equation (3.18), $S_{\text{sw}}$ is also the spectral shape parameter of the bubble, defined as follows:
$S_{\text{sw}}$ $=\left(\frac{f}{f_{\text{sw}}}\right)^{3}\left(\frac{7}{3\left(\frac{f}{f_{\text{sw}}}\right)^{2}+4}\right)^{3.5},$
$f_{\text{sw}}$ $=1.9\times 10^{-5}\left(\frac{1}{v_{w}}\right)\left(\frac{\beta}{H_{*}}\right)\left(\frac{T_{n}}{100}\right)\left(\frac{g_{*}}{100}\right)^{\frac{1}{6}}.$ (3.20)
The density of background GW from turbulence sources is defined as [38]:
@@
$\Omega_{\text{turb}}(f)h^{2}=3.35\times 10^{-4}\left(\frac{\beta}{H_{*}}\right)^{-1}\left(\frac{k_{\text{turb}}\alpha}{1+\alpha}\right)^{\frac{3}{2}}\left(\frac{g_{*}}{100}\right)^{-\frac{1}{3}}v_{w}S_{\text{turb}},$ (3.21)
@@
where $k_{\text{turb}}$ is the magneto-hydrodynamic (MHD) turbulence, defined as follows:
$k_{\text{turb}}=0.05k_{v}$ (3.22)
@@
In equation (3.21), $S_{\text{turb}}$ is defined as follows:
$S_{\text{turb}}$ $=\frac{\left(\frac{f}{f_{\text{turb}}}\right)^{3}}{\left(1+\frac{8\pi f}{h_{*}}\right)\left(1+\frac{f}{f_{\text{turb}}}\right)^{\frac{11}{3}}},\;\;\;h_{*}=1.65\times 10^{-5}\left(\frac{T_{n}}{100}\right)\left(\frac{g_{*}}{100}\right)^{\frac{1}{6}},$
$f_{\text{turb}}$ $=2.7\times 10^{-5}\left(\frac{1}{v_{w}}\right)\left(\frac{\beta}{H_{*}}\right)\left(\frac{T_{n}}{100}\right)\left(\frac{g_{*}}{100}\right)^{\frac{1}{6}}.$ (3.23)

### v3

@@ Abstract
This paper studies gravitational waves in a dark matter model composed of three types of particles with distinct spins, along with a scalar field $\phi$ that mediates interactions between Standard Model particles and dark matter. It discusses the electroweak phase transition following the Big Bang, during which all particles are initially massless due to the inactive Higgs mechanism. As temperature decreases, the effective potential reaches zero at two points, leading to two minima at the critical temperature ($T_{c}$), and eventually to a true vacuum state. The formation of new vacuum bubbles, where electroweak symmetry is broken and particles acquire mass, generates gravitational waves as these bubbles interact with the fabric of space-time. The paper derives the gravitational wave frequency and detection range based on the model’s parameters, aligning with observational data from the Planck satellite and detection thresholds from PandaX-4T and XENONnT for some parameter points. It concludes by comparing the predicted background gravitational wave density with the sensitivities of LISA, BBO and $\mu$-Ares detectors.
@@ 1 Introduction
After the Big Bang, all particles, including dark matter particles undergoes an electroweak phase transition at a certain temperature. We present a dark matter model composed of three types of particles, each with distinct spins, along with a scalar field $\phi$ that serves as an intermediary between Standard Model particles and dark matter, capable of assuming the vacuum expectation value. Prior to electroweak phase transition, three-component dark matter particles and all SM particles are massless, because the Higgs mechanism has not yet activated. Before the electroweak phase transition, the minimum of effective potential at a certain value of the scalar field $\phi$ is equal to zero, which indicates the initial vacuum before the phase transition. As the temperature decreases, the effective potential becomes zero at two distinct points of the scalar field $\phi$, resulting in two minima corresponding to the critical temperature ($T_{c}$). As the temperature continues to drop, the potential reaches a distinct negative value, leading to a single minimum known as the ”true vacuum”. Cosmic phase transitions happen when the temperature decreases below a critical level, resulting in the Universe transitioning from a symmetric phase to one characterized by broken symmetry. This temperature, at which the cores of true vacuum bubbles begin to form in space, is known as the nucleation temperature ($T_{n}$). In fact, with the breaking of symmetry, the scalar field ($\phi$) acquires a nonzero vacuum expectation value (VEV), and all particles gain mass through the Higgs mechanism. New vacuum bubbles form spherically and move through space, replacing the vacuum before the phase transition. Inside the bubbles, electroweak symmetry is broken, and the Higgs mechanism is activated. In fact, inside the bubbles, dark matter and Standard Model particles have mass, whereas outside these bubbles, they are massless [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
@@
The motion of spherically shaped new vacuum bubbles (true vacuum) within the hot plasma resulting from the Big Bang leads to expansion and oscillation in the space-time fabric due to the distribution of mass in space. As bubble walls collide, they induce perturbations in the surrounding space-time. This occurs at the percolation temperature ($T_{p}$) and leads to the emission of gravitational waves that travel through space. At the percolation temperature, bubbles occupy roughly 30% of the Universe’s volume [21]. These gravitational waves were produced in the early universe, and their strength has diminished significantly over time and the expansion of the universe, ultimately transforming into cosmic background gravitational waves. In this paper, we derive the GW frequency and range for the detection of these background gravitational waves, considering the assumption of the existence of multi-component dark matter. The first order electroweak phase transition can be acquired by various dark matter models [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
@@
In this article, we first present the Lagrangian and its dependent constraints for our dark matter model based on WIMPs. In our model, a pair is considered for the Higgs particle, and its mass ($M_{H_{2}}$) is determined by the Gildener–Weinberg mechanism. Next, we select points from our model’s parameter space where their relic density aligns with the Planck satellite report and their rescaled DM-nucleon cross sections exceed the PandaX-4T and XENONnT detection threshold. We then calculate the effective potential for these points and, using the Euclidean action, determine the necessary parameters for background gravitational waves, which we substitute into the relevant equations. Finally, We determine the detection range for the background gravitational wave density using a three-component dark matter model and compare it to the sensitivities of the LISA, BBO and $\mu$-Ares gravitational wave detectors [36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
The paper is organized as follows: Section 2 introduces the Lagrangian of our model, while Section 3 examines the electroweak phase transition via the one-loop effective potential. In Section 4, we present our results, considering the gravitational wave spectrum using nine selected points outlined in Table 2. We conclude in Section 5.
@@ 2.1 Lagrangian of model
dark charge ($Q$) 11 0 0 $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}$
@@
Consequently, the average $H_{2}>$ in the flat direction is calculated as follows:
$H_{2}>=\sqrt{\nu_{1}^{2}+\nu_{2}^{2}}=\nu.$ (2.11)
@@ 2.2 Appropriate points
In this section, we use the micrOMEGAs to select nine points from parameter space of our model that achieve a relic density between 0.96 and 0.14, that some of these points have DM–nucleon cross sections that exceed the detection thresholds of the PandaX-4T and XENONnT detectors [46, 36]. For $M_{DM}>40\;GeV$, the direct detection constraint is expressed as follows:
@@
where $\xi_{S}+\xi_{\psi}+\xi_{V}=1$. Using dark matter relic density measurements from the Planck collaboration, we identify nine suitable points presented in Table 2. In this table, we have obtained the mass of each dark matter component, the vacuum expectation values $(\nu_{2})$ for the ($\phi$) field, the allowed values for the coupling constant $\lambda_{\phi s}$ for the interaction between the scalar field (s) and ($\phi$), the mass of the second Higgs particle (dark Higgs), and the relic density values for the selected points.
Table 2: Characteristics of the nine selected points.
Point $M_{V}$(GeV) $M_{S}$(GeV) $M_{\psi}$(GeV) ${\nu}_{2}$(GeV) $\lambda_{\phi s}$ $M_{H_{2}}$(GeV) ${\Omega}h^{2}$
A 54.5 1125 611 567 8 125.79 0.12
B 330 665 175 340 8 124 0.14
C 300 725 220 400 8 125 0.14
D 800.2 50 600 290 $10^{-4}$ 126.72 0.13
E 1125.25 1500 750.25 630 10 450.65 0.10
F 600.4 430 360.4 430 $10^{-2}$ 120 0.97
G 500.5 287.5 300.5 390 1 84.52 0.96
H 1000.33 533.33 667 630 $10^{-4}$ 203.31 0.97
I 334 50 200.67 310 $10^{-4}$ 34 0.13
@@
The rescaled DM-nucleon cross sections of some selected points, relative to the threshold of the PandaX-4T and XENONnT detectors, are shown in Figure 2.
@@
Figure 2: These figures illustrate the points from table 2 in relation to rescaled DM-Proton cross section versus DM masses, with $\nu_{2}$ and relic density ${\Omega}h^{2}$ represented by color bar.
@@ 3.1 One-loop effective potential
with thermal functions
$J_{\mathbf{B,F}}(x)=\int_{0}^{\infty}dy\,y^{2}\ln\left(1\mp e^{-\sqrt{y^{2}+x^{2}}}\right).$ (3.8)
The $J_{(B_{i})}(x)$ and $J_{(F_{i})}(x)$ are simplified as follows [50]:
@@
$S_{3}(T)=4\pi\int_{0}^{\infty}dr\,r^{2}\left(\frac{1}{2}\left(\frac{d\varphi}{dr}\right)^{2}+V_{\text{eff}}(\varphi,T)\right),$ (3.12)
@@
The nucleation temperature ($T_{n}$) and percolation temperature ($T_{p}$) occur where the corresponding euclidean action are $S_{n}=\frac{S_{3}(T_{n})}{T_{n}}\simeq 140$ [39, 35] and $S_{p}=\frac{S_{3}(T_{p})}{T_{p}}\simeq 100$ [42], respectively. To solve equation (3.13) and find the Euclidean action (3.12), we used the AnyBubble package [38]. This package allows us to calculate the Euclidean action for each selected point using the two minima in the effective potential diagram as function of the scalar field ($\phi$), and plot the ratio of the Euclidean action to temperature (T). In figure 3, we present $\frac{S_{3}(T)}{T}$ for selected points in table 2.
@@
Figure 3: The blue line shows $S_{3}(T)/T$ as a function of T, while the dashed red and purple lines indicate the values $S_{3}(T)/T=140$ and $S_{3}(T)/T=100$, respectively, at which nucleation and percolation occur. The black dots are derived using the AnyBubble package.
@@
In Figure 3, the nucleation temperature ($T_{n}$) and percolation temperature ($T_{p}$) are determined by finding the intersection points of the curve fitting and the horizontal lines. To determine the critical temperature $T_{c}$ for the points in table 2, we must plot the effective potential (3.1) against the scalar field ($\phi$) to identify the temperature at which the effective potential exhibits two aligned minima (see figure 4).
@@
Figure 4: Effective potential diagrams of points in table 3 as a function of $\phi$ scalar field at percolation ,nucleation, critical, and supercritical temperatures.
@@
Figure 4 shows the effective potential for each point in Table 2 at the four temperatures: percolation ($T_{p}$), nucleation ($T_{n}$), critical ($T_{c}$), and supercritical ($T>T_{c}$). To ensure $V_{\text{eff}}(0,T)=0$ at all temperatures, we subtract a constant term from the potential: $V_{\text{eff}}(\phi,T)\rightarrow V_{\text{eff}}(\phi,T)-V_{\text{eff}}(0,T)$. As shown in Figure 4, at the nucleation temperature ($T_{n}$), the zero minimum is known as a false vacuum and the negative minimum is known as a true vacuum. There is a bulge between the two minima, which prevents the electroweak phase transition from occurring smoothly between the minima, and instead occurs suddenly and explosively, like a dam breaking.
@@
The supercritical temperatures ($T>T_{c}$) shown in Figure 4 are determined as arbitrary and indicate the position of the effective potential in relation to the percolation, nucleation and critical temperatures. Figure 4 illustrates that as the temperature decreased, the effective potential also diminishes. Also, for all selected points, the thermal order is as follows:
$T_{c}>T_{n}>T_{p}.$ (3.15)
As the separation between the critical and nucleation temperatures grows, more vacuum energy is released during the false-to-true vacuum transition. This regime is known as supercooling. The released energy can temporarily heat the primordial plasma, a process called reheating, but it is not enough to raise the negative minimum in Figure 4 above the zero minimum and to create a reverse phase transition. During reheating, the higher temperature suppresses the formation of new bubbles, effectively increasing the spacing between bubble nuclei and allowing existing bubbles to grow larger before colliding. Larger bubble collisions produce stronger gravitational waves. Thus, supercooling enhances reheating, which in turn strengthens the gravitational-wave signal.
The difference between the critical and nucleation temperatures determines the degree of supercooling. Therefore, the supercooling parameter is as follows:
$\delta_{sc,n}=\frac{T_{c}-T_{n}}{T_{c}}.$ (3.16)
When bubbles formation occurs along with supercooling, the released heat reaches a level where it can heat significant the initial plasma. As the system cools, the plasma temperature returns to the nucleation temperature. This time delay gives true vacuum bubbles enough time to occupy about 30% of the universe’s volume. Therefore, in the temperature versus $\delta_{sc,n}$ plot, for $\delta_{sc,n}\gtrsim 0.5$, the percolation-temperature curve tends to a constant value and no longer decreases [21]. In Table 3, the supercooling parameter $\delta_{sc,n}$ is calculated for each selected points.
Table 3: Thermal characteristics of the selected points in Table 2.
Point $T_{p}$(GeV) $T_{n}$(GeV) $T_{C}$(GeV) ${\phi}_{V_{min}}]_{T_{p}}$(GeV) ${\phi}_{V_{min}}]_{T_{n}}$(GeV) $\delta_{sc,n}$
A 112 116 162.5 607.471 606.022 0.28
B 98.5 101 128.2 404.004 402.363 0.21
C 104.7 106.7 135.3 450.592 449.134 0.21
D 120 126 174.5 396.343 397.913 0.27
E 338 350 396 665.93 662.441 0.11
F 96 98 132 486.955 486.125 0.25
G 68 70 99 455.481 454.619 0.29
H 160 166 222.5 671.402 670.249 0.25
I 31.5 32 51 395.154 395.05 0.37
@@
In Table 3, the ${\phi}_{V_{min}}]_{T_{n}}$ denotes the point of $\phi$} scalar field at which the effective potential value in $T_{n}$ is minimized, as illustrated in Figure 4. In this context, the ${\phi}_{V_{min}}]_{T_{n}}$ represents the true vacuum. According to the values in the $\delta_{sc,n}$ column of Table 3, supercooling ($\delta_{sc,n}\gtrsim 0.5$) has not occurred for the selected points in our model.
@@ 3.2 Gravitational waves
$\alpha=\frac{\Delta\left(V_{\text{eff}}-T\frac{\partial V_{\text{eff}}}{\partial T}\right)_{\big|_{T_{p}}}}{\rho_{*}},\;\;\;\rho_{*}=\frac{\pi^{2}g_{*}}{30}T_{p}^{4}$
$\frac{\beta}{H_{*}}=T_{p}\left.\frac{d}{dT}\left(\frac{S_{3}(T)}{T}\right)\right|_{T_{p}},$ (3.17)
@@
where $g_{*}=100$. In equation (3.2), we obtain the parameter $\frac{\beta}{H_{*}}$ from the slope of the graphs in figure 3 at the intersection point of the curve fitting and the purple line. Now, we derive the density relations of gravitational waves for all three types of their sources, including collision (coll), sound wave (sw), and turbulence (turb) [52]. We begin with the density of background gravitational waves from collision sources [53, 54]:
@@
$\Omega_{\text{coll}}^{env}(f)h^{2}=\quad 1.67\times 10^{-5}\,\Delta\left(\frac{100}{g_{*}}\right)^{\frac{1}{3}}\left(\frac{H_{*}}{\beta}\right)^{2}\left(\frac{\kappa_{\phi}\alpha}{1+\alpha}\right)^{2}S_{\mathrm{env}}(f/f_{\mathrm{env}}),$ (3.18)
with
$\Delta=\frac{0.48v_{w}^{3}}{1+5.3v_{w}^{2}+5v_{w}^{4}},$ (3.19)
@@
$S_{\mathrm{env}}(r)$ is the spectral shape parameter of the bubble, defined by the following relation:
$S_{\mathrm{env}}(r)=\left(0.064r^{-3}+0.456r^{-1}+0.48r\right)^{-1},$ (3.20)
where $r={f}/{f_{env}}$, and $f$ is the GW frequency of the gravitational wave, with peak frequency at:
$\frac{f_{\mathrm{env}}}{1\,\mathrm{\mu Hz}}=16.5\left(\frac{f_{*}}{\beta}\right)\left(\frac{g_{*}}{100}\right)^{\frac{1}{6}}\left(\frac{\beta}{H_{*}}\right)\left(\frac{T_{*}}{100\,\mathrm{GeV}}\right),$
$\frac{f_{*}}{\beta}=\frac{0.35}{1+0.069v_{w}+0.69v_{w}^{4}}.$ (3.21)
In equation (3.18), $\kappa_{\phi}$ is the fraction of latent heat converted into the gradient energy of the Higgs-like field, defined as follows [55]:
$\kappa_{\phi}=\frac{1}{1+0.715\alpha}\left[0.715\alpha+\frac{4}{27}\sqrt{\frac{3\alpha}{2}}\right],$ (3.22)
and $v_{w}$ is the bubble propagation speed [56].
The density of background GW from sound wave sources is defined as [57, 58]:
$\Omega_{\text{sw}}(f)h^{2}=2.061\,F_{gw,0}\,\Gamma^{2}\,\bar{U}_{f}^{4}\,S_{sw}(f)\,\tilde{\Omega}_{gw}\min\left(H_{*}R_{*}/\bar{U}_{f},1\right)\left(H_{*}R_{*}\right)h^{2},$ (3.23)
where
$F_{gw,0}=3.57\times 10^{-5}\left(\frac{100}{g_{*}}\right)^{\frac{1}{3}},$
$R_{*}=\left(8\pi\right)^{1/3}\frac{v_{w}}{\beta},$
$K=\frac{\kappa_{sw}\alpha}{1+\alpha}=\Gamma\bar{U}_{f}^{2},$ (3.24)
the $R_{*}$ and $K$ are bubble separation and kinetic energy fraction, respectively.
In equation (3.23), the $\Gamma={4}/{3}$ and $\tilde{\Omega}_{gw}\sim 0.012$ are ratio of enthalpy to the energy density and determined from simulations, respectively. Also, for fluid the spectral shape parameter $S_{\text{sw}}(f)$ of the bubble, defined as follows [59, 42]:
$S_{sw}(f)$ $=\left(\frac{f}{f_{sw}}\right)^{3}\left(\frac{7}{4+3(f/f_{sw})^{2}}\right)^{\frac{7}{2}},$
$\frac{f_{sw}}{1\,\mu\mathrm{Hz}}$ $=2.6\left(\frac{z_{p}}{10}\right)\left(\frac{T_{*}}{100\,\mathrm{GeV}}\right)\left(\frac{g_{*}}{100}\right)^{\frac{1}{6}}\left(\frac{1}{H_{*}R_{*}}\right),\;\;\;z_{p}\sim 10.$ (3.25)
the $\kappa_{sw}$ in equation (3.24) is determined as [60]:
$\kappa_{sw}=\begin{cases}\displaystyle\frac{c_{s}^{11/5}\kappa_{A}\kappa_{B}}{(c_{s}^{11/5}-v_{w}^{11/5})\kappa_{B}+{v_{w}}{c_{s}}^{6/5}\kappa_{A}},&v_{w}\leq c_{s}\\ \kappa_{B}+(v_{w}-c_{s})\delta\kappa+\frac{(v_{w}-c_{s})^{3}}{(v_{J}-c_{s})^{3}}l_{k},&c_{s}<v_{w}<v_{J}\\ \displaystyle\frac{(v_{J}-1)^{3}v_{J}^{5/2}v_{w}^{-5/2}\kappa_{C}\kappa_{D}}{[(v_{J}-1)^{3}-(v_{w}-1)^{3}]v_{J}^{5/2}\kappa_{C}+(v_{w}-1)^{3}\kappa_{D}},&v_{J}\leq v_{w}\end{cases}$ (3.26)
where
$\kappa_{A}$ $\simeq v_{w}^{6/5}\frac{6.9\alpha}{1.36-0.037\sqrt{\alpha}+\alpha},$
$\kappa_{B}$ $\simeq\frac{\alpha^{2/5}}{0.017+(0.997+\alpha)^{2/5}},$
$\kappa_{C}$ $\simeq\frac{\sqrt{\alpha}}{0.135+\sqrt{0.98+\alpha}},$
$\kappa_{D}$ $\simeq\frac{\alpha}{0.73+0.083\sqrt{\alpha}+\alpha},$
$\delta\kappa$ $\simeq-0.9\log\frac{\sqrt{\alpha}}{1+\sqrt{\alpha}},$
$l_{\kappa}$ $\simeq\kappa_{C}-\kappa_{B}-(v_{J}-c_{s})\delta\kappa,$
$v_{J}$ $=\frac{1}{1+\alpha}\left(c_{s}+\sqrt{\alpha^{2}+\frac{2\alpha}{3}}\right).$ (3.27)
The density of background GW from turbulence sources is defined as [61, 62]:
@@
$\Omega_{\text{turb}}(f)h^{2}=3.35\times 10^{-4}v_{w}\left(\frac{H_{*}}{\beta}\right)\left(\frac{\kappa_{\text{turb}}\alpha}{1+\alpha}\right)^{3/2}\left(\frac{100}{g_{*}}\right)^{1/3}\frac{r^{3}}{(1+r)^{11/3}(1+8\pi f/H_{0})}$ (3.28)
@@
where $r={f}/{f_{turb}}$ and $\kappa_{turb}$ is the magneto-hydrodynamic (MHD) turbulence, defined as follows [57]:
$\kappa_{turb}\simeq(0.05\;\;to\;\;0.1)\times\kappa_{sw}.$ (3.29)
@@
In equation (3.28), the peak frequency $f_{\text{turb}}$ and the red-shifted Hubble rate at GW generation $H_{0}$ are defined as follows:
$f_{\text{turb}}={27}\frac{1}{v_{w}}\left(\frac{\beta}{H_{n}}\right)\left(\frac{T_{*}}{100\,\text{GeV}}\right)\left(\frac{g_{*}}{100}\right)^{1/6}\mu\text{Hz},$
$H_{0}=16.5\left(\frac{g_{*}}{100}\right)^{\frac{1}{6}}\left(\frac{T_{*}}{100\,\text{GeV}}\right)\,\mu\text{Hz}.$ (3.30)